OpenStudy (anonymous):

s=ut=0.5at^2 a=9.8 t=5 u=7

OpenStudy (bibby):

Let's start on the right hand side.\[\huge 0.5at^2=0.5*a*(t^2)\]

OpenStudy (anonymous):

122.5 is the answer for ^

OpenStudy (anonymous):

so is it 12+122.5

OpenStudy (anonymous):

or 35+122.5

OpenStudy (bibby):

t^2 = 49 49*0.5*9.8 = ?

OpenStudy (bibby):

Oh wait, why are there 2 equal signs in the equation?

OpenStudy (anonymous):

it is meant to be a +

OpenStudy (bibby):

Can you rewrite the whole thing then?

OpenStudy (anonymous):

so is it 12+122.5 or 35+122.5

OpenStudy (anonymous):

s=ut+0.5at^2 a=9.8 t=5 u=7

OpenStudy (anonymous):

^^^^^^^^

OpenStudy (bibby):

ah ok. so we know ut = ?

OpenStudy (anonymous):

uhhhh, i think 35 or 12

OpenStudy (anonymous):

but i am 80% on 35 and 20% on 12

OpenStudy (bibby):

u*t. plug in the variables u = 7 t = 5

OpenStudy (anonymous):

so it is 35

OpenStudy (bibby):

lol, pls don't revisit that. That is my name. correct, 35

OpenStudy (anonymous):

okay, so it is 35+122.4=157.4

OpenStudy (bibby):

how are you getting 122.4? 0.5*9.8*(7^2)

OpenStudy (anonymous):

240.1

OpenStudy (bibby):

it's important to note the parentheses. \[\large 0.5at^2=0.5*9.8*(t^2)\]

OpenStudy (bibby):

yep

OpenStudy (anonymous):

so 34+240.1=274.1

OpenStudy (bibby):

35* s= 35+240.1 = 275.1

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

I suggest you change your description and take the ask.fm link

OpenStudy (anonymous):

LOL

OpenStudy (bibby):

lol I must keep it there for the greater good. The people want to ask me things

OpenStudy (anonymous):

Yeah i agree

OpenStudy (anonymous):

but you can see pictures like the V

OpenStudy (bibby):

That's not a V. It's a camera inside of a mouth. That's the throat.

OpenStudy (anonymous):

Yes that is what i mean the throat i call it the V the thing you mentioned i call the veg

OpenStudy (bibby):

Exactly. Rest assured, that isn't a picture of female body parts.

OpenStudy (anonymous):

Yeah i know

OpenStudy (bibby):

do we have more math to do?

OpenStudy (anonymous):

No i finished Aviv

OpenStudy (bibby):

Excellent, Jiggly puff

OpenStudy (anonymous):

Call me Artix it is on my profile page

OpenStudy (anonymous):

it is my real name

OpenStudy (bibby):

Isn't that the guy that made adventure quest's name?

OpenStudy (anonymous):

Really

OpenStudy (anonymous):

I have never heard of adventure quest what is it?

OpenStudy (bibby):

a game https://twitter.com/ArtixKrieger

OpenStudy (anonymous):

????????? He has the same name except mine is Kreiger but his is Krieger

OpenStudy (anonymous):

Wow he is on so many games is he a kind of boss on games

OpenStudy (anonymous):

Okay that is freaky same name

OpenStudy (bibby):

Exactly. That sounds like a bit more than a coincidence. Either way, pleasure meeting you artix

OpenStudy (anonymous):

Pleasure meeting you today too Aviv

OpenStudy (anonymous):

Don't know

OpenStudy (anonymous):

Okay i am going to leave go to go to sleep it is 9

OpenStudy (anonymous):

bye gotta log out

OpenStudy (bibby):

gnight